NEW HIGH-RESOLUTION ABSORPTION CROSS-SECTION MEASUREMENTS OF HCFC-142B IN THE MID-IR

\underline{KARINE LE BRIS}, Department of Physics, St. Francis Xavier University, Antigonish, NS, Canada, B2G2W5; KIMBERLY STRONG, Department of Physics, University of Toronto, Toronto, ON, Canada, M5S1A7; and STELLA MELO, Canadian Space Agency, St. Hubert, QC, Canada, J3Y8Y9.

HCFC-142b (1-chloro-1,1-difluoroethane) is a temporary substitute for ozone-depleting chlorofluorocarbons (CFCs). However, due to its high absorption cross-sections in the mid-IR, HCFC-142b is also a highly potent greenhouse gas, now detectable from space by satellite missions. So far, the accuracy of the retrieval has been limited by the lack of reference data in a range of temperatures compatible with atmospheric observations. We present new absorption cross section measurements of HCFC-142b at high-resolution (0.02 cm^{−1}) from 223 K to 283 K in the 600 cm^{−1}– 4000 cm^{−1} spectral window. The composite spectra are calculated for each temperature from a set of acquisitions at different pressures by Fourier transform spectroscopy.